OpenStudy (pokemon23):

Sam has $2.05 in quarters and dimes. He has 4 more quarters than dimes.Find the number he has of each coin. Sam has $2.05 in quarters and dimes. He has 4 more quarters than dimes.Find the number he has of each coin. @Mathematics

OpenStudy (pokemon23):

i need an equation?

OpenStudy (pokemon23):

quarters are .25cents

OpenStudy (pokemon23):

dimes are .10cents

OpenStudy (lilg132):

y = (0.25 * 4x) + (0.1x)

OpenStudy (anonymous):

But there are two variables.

OpenStudy (pokemon23):

$2.05= (0.25 * 4x) + (0.1x) is't this correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

please give me a medal

OpenStudy (pokemon23):

guide me

OpenStudy (anonymous):

ok just a minute.

OpenStudy (lilg132):

lol i wrote the equation but i get no medal :p

OpenStudy (lilg132):

ok well the way to solve it is first do 0.25 * 4 = 1 then do 1 + 0.1 = 1.1 then 2.20 / 1.1 = 2 which means there were 2 dimes and four times as many dimes were quarters which means there were 8 quarters

OpenStudy (anonymous):

yes correct.

OpenStudy (lilg132):

did you understand this pokemon?

OpenStudy (pokemon23):

you forgot the variable?

OpenStudy (lilg132):

x = 2

OpenStudy (anonymous):

which variable?

OpenStudy (pokemon23):

x variable

OpenStudy (anonymous):

I'm very confused about dimes and quarters, if you need help in solving the equation, I can help you.

OpenStudy (lilg132):

when you do 2.2 / 1.1 = 2 you have worked out what x is

OpenStudy (anonymous):

x = 2

OpenStudy (lilg132):

dimes = 10 cents quarters = 25 cents

OpenStudy (pokemon23):

a dime is 10 cents and quarters are 25 cents

OpenStudy (anonymous):

and how many cents make a dollar?

OpenStudy (lilg132):

100

OpenStudy (lilg132):

lol

OpenStudy (lilg132):

im not even from usa i know that :p

OpenStudy (pokemon23):

isn't suppose to be 2.05/1.1

OpenStudy (anonymous):

Hey, Sam has four MORE quarters not four TIMES.

OpenStudy (lilg132):

no four times more quarters

OpenStudy (pokemon23):

then this equation is wrong?

OpenStudy (lilg132):

and pokemon23 did you make up this question about $2.05 because it doesnt seem possible to me with just dimes and quarters

OpenStudy (anonymous):

I think so.

OpenStudy (pokemon23):

nah its from my math textbook

OpenStudy (lilg132):

oh wait your right you changed the wording to 4 more instead of 4 times

OpenStudy (lilg132):

so it will be a different equation

OpenStudy (anonymous):

Yes, you would need a whole lot of dimes and quarters.

OpenStudy (pokemon23):

can you guys give me the correct equation ? please

OpenStudy (anonymous):

I still think you can't have $2.05 with just dimes an quarters. I mean you can, but you'll need a whole bag of them.

OpenStudy (pokemon23):

?

OpenStudy (lilg132):

ok let dimes = x quarters = x + 4 0.25(x+4) + 0.1x = 2.05 0.25x + 1 + 0.1x = 2.05 0.25x + 0.1x = 2.05 - 1 0.35x = 1.05 x = 3 3 dimes and 7 quarters = $2.05

OpenStudy (pokemon23):

ty